A speed skater goes around a turn that has a radius of 27 m. The skater has a speed of 15 m/s and experiences a centripetal force of 470 N. What is the mass of the skater?
Answer #1

centripetal force = mv^2/r

mv^2/r = 470 N

m = 470*r/v^2

m = 470*27/15^2 = 56.4 kg

MAss of skater = 56.4 kg
Answer #2

use this
Fc = m v^2 / r
470 = m (15^2) / 27
m = 56.42 kg